OpenStudy (anonymous):

BEST ANSWER THE QUESTION GETS THE ANSWER!!!!! Simplify the following expression: (5x^2 + 3x + 4) − (2x^2 + 5x − 1). If the final answer is written in the form Ax^2 + Bx + C, what is the value of A? Numerical Answers Expected!

OpenStudy (mertsj):

3

OpenStudy (whpalmer4):

\[(5x^2+3x+4)-(2x^2+5x-1) = (5x^2+3x+4)-1(2x^2+5x-1)\]Distribute the -1 on the second part\[(5x^2+3x+4)-2x^2-5x-(-1)=5x^2+3x+4-2x^2-5x+1\]Collect like terms\[5x^2-2x^2+3x-5x+4+5 \]\[= 3x^2-2x+9\]Compare with \[Ax^2+Bx+C\]